OpenStudy (anonymous):

A new oil filter costs 180 and oil costs 134 per liter . Write equation for the C of an oil changes as a function of the number of liters of oil used (u): If Mr.Aquino's car requires 3.4 liters of oil , how much would change oil cost Mr. Aquino?

OpenStudy (zzr0ck3r):

so we are going to have to pay 180 no matter what. then we have to pay for each liter knowing the form will be like this f(x) = mx+b what do you think m is and what do you think b is?

OpenStudy (anonymous):

m is 134 and b is 3.4 ?

OpenStudy (anonymous):

then the answer would be 455.6 ? am i correct?

OpenStudy (zzr0ck3r):

we are going to pay 180 no matter what f(x) = 180 then we are going to pay for each liter at 134 so f(x) = 180 + 134x where x is the amount of liters.

OpenStudy (zzr0ck3r):

now we have the equation, now we plug in 3.4 for x to find f(3.4)

OpenStudy (zzr0ck3r):

if you dont understand, please ask a question.

OpenStudy (anonymous):

f(3.4)=180+134x

OpenStudy (zzr0ck3r):

put a 3.4 where you have an x

OpenStudy (anonymous):

f(3.4)=180+134(3.4)

OpenStudy (zzr0ck3r):

\(f(x) = 180+134x\) so what ever we plug in for \(x\) we plug in for \(x\) everywhere. \(f(2) = 180+134*2\\f(34343.43)=180+134*34343.43\\f(\checkmark)=180+134*\checkmark\\f(3.4) = 180+134*3.4\)

OpenStudy (anonymous):

I don't get it . I'm not good in math.

OpenStudy (anonymous):

@Sakura22 This is just like the previous question, C=180+134u. Where, C is the total cost and u is the number of liters of oil used. So, it is given Mr.Aquino's car consumes 3.4 liters of oil, Here, 3.4 is u and you have to find C. So, C=180+(134*3.4) C= 180+455.6= 635.6 Understood?

OpenStudy (anonymous):

yeah. thank you again . :)